# M1: Towards Scalable Test-Time Compute with Mamba Reasoning Models

### v1

@@ Abstract
Effective reasoning is crucial to solving complex mathematical problems. Recent large language models (LLMs) have boosted performance by scaling test-time computation through long chain-of-thought reasoning. However, transformer-based models are inherently limited in extending context length due to their quadratic computational complexity and linear memory requirements. In this paper, we introduce a novel hybrid linear RNN reasoning model, M1, built on the Mamba architecture, which allows memory-efficient inference. Our approach leverages a distillation process from existing reasoning models and is further enhanced through RL training. Experimental results on the AIME and MATH benchmarks show that M1 not only outperforms previous linear RNN models but also matches the performance of state-of-the-art Deepseek R1 distilled reasoning models at a similar scale. We also compare our generation speed with a highly performant general purpose inference engine, vLLM, and observe more than a 3x speedup compared to a same size transformer. With throughput speedup, we are able to achieve higher accuracy compared to DeepSeek R1 distilled transformer reasoning models under a fixed generation time budget using self-consistency voting. Overall, we introduce a hybrid Mamba reasoning model and provide a more effective approach to scaling test-time generation using self-consistency or long chain of thought reasoning.
@@ 1 Introduction
Overall, we show that M1 performs on par with DeepSeek-R1-Distill-Qwen-1.5B, achieving scores of 82 on MATH500 (Hendrycks et al., 2021), 22 on AIME25 (MAA, 2025), 23 on AIME24 (MAA, 2024), and 44 on OlympiadBench (He et al., 2024), while offering 3x faster inference throughput, even compared to the highly optimized vLLM (Kwon et al., 2023) implementation for Transformer models.
@@ 2.2 Enhancing Reasoning via Scaled Inference Compute
Increasing the computational budget during inference has become a promising approach to boost LLM performance. Methods like Chain of Thought (CoT) and its derivatives have achieved notable gains on reasoning benchmarks by breaking down complex tasks into intermediate steps (Wei et al., 2023; Yao et al., 2023). Although decomposing tasks improves reasoning, it also lengthens generation sequences and raises computational costs. Some recent studies even indicate that this extra computation might itself enhance model capabilities (Pfau et al., 2024). In addition, adaptive compute allocation during inference has been explored. For example, Goyal et al. (2024) incorporated pause tokens into the vocabulary, allowing models to distribute compute more efficiently and improve both reasoning and overall task performance.
@@ 2.3 Alternatives to Transformer Architectures
Even though most reasoning models are based on the Transformer architecture (Grattafiori et al., 2024; Qwen et al., 2025), alternatives have been proposed to alleviate their high computational cost. Models built on top of RNNs (Beck et al., 2024; Peng et al., 2023), state space models (SSMs) (Gu et al., 2022; Gu & Dao, 2024), and linear attention mechanisms (Katharopoulos et al., 2020; Yang et al., 2024) demonstrate superior inference and memory efficiency, particularly for long-context tasks and large-batch generation. The Mamba series (Mamba-1 and Mamba-2) notably introduced selective state spaces to enable linear-time sequence modeling with strong performance (Gu & Dao, 2024; Dao & Gu, 2024). In addition, hybrid architectures that combine a few self-attention layers with subquadratic layers (e.g., Mamba) have emerged, showing advantages over both pure Transformer and pure subquadratic designs (Lieber et al., 2024; Ren et al., 2024; Dong et al., 2024). Such architectures are particularly suited to meet the high compute demands of inference-time scaling, and our work investigates their scaling properties.
@@ 2.4 Knowledge Distillation Strategies
Knowledge distillation has proven to be an effective means of transferring capabilities from large teacher models to smaller, more efficient student models (Hinton et al., 2015). In LLMs, this process compresses a larger pre-trained model into a more compact version while preserving core knowledge and functionality (Gu et al., 2024; Xu et al., 2024). Although larger models tend to exhibit superior reasoning abilities due to scaling properties (Xu et al., 2025; Wei et al., 2022), distillation techniques have enabled smaller models to achieve competitive reasoning performance (DeepSeek-AI et al., 2025; Labs, 2025). While most efforts have focused on intra-architecture distillation (e.g., Transformer-to-Transformer), recent studies have ventured into cross-architecture distillation. For instance, pretrained Transformers have been distilled into architectures such as RNNs (Kasai et al., 2021; Mercat et al., 2024), linear attention models (Zhang et al., 2024b; ), convolutional networks (Ralambomihanta et al., 2024), and SSMs (Bick et al., 2024; Wang et al., 2024; Paliotta et al., 2025). Whether the robust reasoning abilities of Deepseek R1 (DeepSeek-AI et al., 2025) distilled models can be effectively transferred across different architectures remains an open question.
@@ 3 The M1 Reasoning Model
In this section, we present a multi-stage process for building our hybrid linear RNN reasoning model, M1. The approach has three stages: distillation, SFT, and RL. We begin by distilling a Transformer model into a Mamba architecture, adapting the method of Wang et al. (2025), which initializes the hybrid model’s weights from a transformer model. We then perform math-specific supervised fine-tuning (SFT) on general mathematical datasets to enhance the model’s mathematical performance, first without yet incorporating datasets generated by reasoning-focused models, and then with reasoning data leveraging multiple large-scale datasets generated by the R1 model series. Finally, we apply R1’s GRPO method to further enhance the model’s math reasoning capability.
@@
The first step in building our M1 model is distilling a pretrained transformer model into a Mamba model. We adapt the distillation approach introduced by  Wang et al. (2025).
[…]
@@
9:     $\ {\mathbf{B}}_{t}:(B,N)\leftarrow{\mathbf{W}}^{K}{\bm{o}}_{t}$
10:     $\ {\mathbf{C}}_{t}:(B,N)\leftarrow{\mathbf{W}}^{Q}{\bm{o}}_{t}$
@@
The MambaInLlama framework (Wang et al., 2025) proposes distilling hybrid Transformer-Mamba models by reusing weights from attention layers. In this distillation procedure, outlined in Algorithm 1, linear projections for ${\mathbf{Q}}$, ${\mathbf{K}}$, ${\mathbf{V}}$, and ${\mathbf{O}}$ are initialized from the corresponding projections for ${\mathbf{C}}$, ${\mathbf{B}}$, ${\mathbf{X}}$, and ${\mathbf{O}}$, respectively. The newly introduced parameters in the Mamba layers are the sampling rate $\Delta$ and the dynamic parameter ${\mathbf{A}}$, which control the resulting Mamba module via a discretization function. Specifically, the sampling rate $\Delta\in\mathbb{R}^{N^{\prime}}$ discretizes ${\mathbf{B}}_{t},{\mathbf{C}}_{t}\in\mathbb{R}^{N\times 1}$, yielding $\overline{{\mathbf{B}}}_{t},\overline{{\mathbf{C}}}_{t}\in\mathbb{R}^{N^{\prime}\times N\times 1}$, as detailed in Algorithm 1. Different from Wang et al. (2025), we introduce two additional linear layers to project from head_dim * kv_head to head_dim * n_head. This is because GQA  (Ainslie et al., 2023) is used in the transformer model to reduce the KV cache. As Mamba does not utilize a KV cache, this expansion can increase the expressiveness of $\mathbf{B}$ and $\mathbf{X}$.
@@
After the initial fine-tuning stage, we finetune on an additional set of math problems and solutions generated by reasoning models. We collect a mixed reasoning dataset, including OpenR1-Math-220k  [^2] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k, OpenThoughts-114k-math [^3] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k, and ServiceNow-AI-R1-Distill [^4] https://huggingface.co/datasets/ServiceNow-AI/R1-Distill-SFT, Magpie-Reasoning-250K [^5] https://huggingface.co/datasets/Magpie-Align/Magpie-Reasoning-V2-250K-CoT-Deepseek-R1-Llama-70B for a total of 8B reasoning tokens. The first two datasets were generated from R1, while the last two was generated from the R1 distilled Qwen 32B model and R1 distilled Llama 70B model. We extended the training length to 24,576 because we found that it covers 99% of the data items. We train the model for five epochs using the same optimizer as before but changing the peak learning rate to $6\times 10^{-6}$.
@@ 4 Experiments
[…]
@@ 4.1 Reasoning Evaluation
Model AIME25 AIME24 MATH500 OlympiadBench AMC23
Pass@1 Maj@32 Pass@1 Maj@32 Pass@1 Maj@32 Pass@1 Pass@1
GPT-4o-0513 9.3 - - - 74.6 - - -
Claude-3.5-1022 16.0 - - - 78.3 - - -
Qwen2.5-Math-7B-Ins - - 13.3 - 79.8 - 40.7 50.6
R1-Llama-8B 32.9 - 50.4 - 89.8 - 54.2 -
R1-Qwen-1.5B 23.5 - 28.9 83.9 91.0 43.3 62.9
M1-3B 22.0 30.8 23.0 45.0 81.7 89.5 43.6 56.0
Table 1: Results for M1-3B, DeepSeek-R1-Distill-Qwen-1.5B, and Qwen2.5-Math-7B-Instruct.
@@
We report the accuracy of M1-3B and DeepSeek-R1-Distill-Qwen-1.5B in Table 1. We use the baseline DeepSeek-R1-Distill-Qwen-1.5B since a 3B R1 reasoning model is still not available. Although M1-3B has more parameters than DeepSeek-R1-Distill-Qwen-1.5B, its speed is still comparable even with shorter contexts, so we believe this is a fair comparison. Our model’s performance is competitive with state-of-the-art open reasoning models in the same model size range and outperforms larger nonreasoning math transformer models. Our model performs slightly worse on AIME24 compared to the DeepSeek-R1-Distill-Qwen-1.5B model. Notably, DeepSeek-R1-Distill-Qwen-1.5B is built on top of the Qwen2.5 MATH models, which were finetuned with over 1T MATH tokens on top of the Qwen2.5 models, significantly more training data than what M1-3B used in total.
@@ 4.3 Test-Time Scaling
The left side of Figure 3 shows the effect of scaling the number of generated samples (while fixing the maximum decoding length) on AIME24 accuracy. Both the baseline model and M1 see increasing accuracy as the number of samples increases, with M1 nearly matching the baseline performance for larger sample sizes. The efficient generation of M1 also means that generating large number of samples at test-time is faster than for the baseline transformer model.
@@ 5 Analysis
Distill + SFT(MATH) + SFT(Reason) 74 17
Distill + SFT(MATH) + SFT(Reason) + RL 82 23
@@
With more efficient models, we can increase the length of sequences used in RL training, resulting in improved performance. Empirically, we see this in Figure 5, which shows an increase in accuracy on AIME25 as we scale up the length of sequences generated when training with GRPO. Training with sequences of maximum length 4096 results in accuracy below 7.5%, while allowing sequences up to length 24K boosts the accuracy up to 22.5%.
@@
Direct Distillation from Reasoning Models We also attempted to distill from Deepseek-R1-Qwen-1.5B instead of Llama-3.2-3B. In this case, we did not SFT on OpenMathInstruct, and instead only SFT on the 8B reasoning data that we collected after distillation. We found that the distilled model’s performance was poor (38% and 3.3% pass@1 accuracy on MATH500 and AIME24, resspectively). Our hypothesis for why this occurs is that 8B tokens is insufficient to effectively transfer reasoning skills from the transformer to Mamba. Although curating a high-quality reasoning dataset demands significant time and effort, we begin by leveraging the standard MATH distillation dataset from OpenMathInstruct (Toshniwal et al., 2024) to first distill a strong MATH model. We then transform this MATH model into a reasoning model via SFT on the dedicated reasoning dataset. This approach achieves strong performance with a much smaller number of reasoning tokens.

### v3

@@ Abstract
Effective reasoning is crucial to solving complex mathematical problems. Recent large language models (LLMs) have boosted performance by scaling test-time computation through long chain-of-thought reasoning. However, transformer-based models are inherently limited in extending context length due to their quadratic computational complexity and linear memory requirements. In this paper, we introduce a novel hybrid linear RNN reasoning model, M1, built on the Mamba architecture, which allows memory-efficient inference. Our approach leverages a distillation process from existing reasoning models and is further enhanced through RL training. Experimental results on the AIME and MATH benchmarks show that M1 not only outperforms previous linear RNN models but also matches the performance of state-of-the-art Deepseek R1 distilled reasoning models at a similar scale. We also compare our generation speed with a highly performant general purpose inference engine, vLLM, and observe more than a 3x speedup compared to a same size transformer. With throughput speedup, we are able to achieve higher accuracy compared to DeepSeek R1 distilled transformer reasoning models under a fixed generation time budget using self-consistency voting. Overall, we introduce a hybrid Mamba reasoning model and provide a more effective approach to scaling test-time generation using self-consistency or long chain of thought reasoning. Code and pre-trained checkpoints are open-sourced at github.com/jxiw/M1.
@@ 1 Introduction
Overall, we show that M1 performs on par with DeepSeek-R1-Distill-Qwen-1.5B, achieving scores of 82 on MATH500 (Hendrycks et al., 2021), 23 on AIME25 (MAA, 2025), 28 on AIME24 (MAA, 2024), and 47 on OlympiadBench (He et al., 2024), while offering 3x faster inference throughput, even compared to the highly optimized vLLM (Kwon et al., 2023) implementation for Transformer models.
@@ 2.2 Enhancing Reasoning via Scaled Inference Compute
Increasing the computational budget during inference has become a promising approach to boost LLM performance. Methods like Chain of Thought (CoT) and its derivatives have achieved notable gains on reasoning benchmarks by breaking down complex tasks into intermediate steps (Wei et al., 2023; Yao et al., 2023). Although decomposing tasks improves reasoning, it also lengthens generation sequences and raises computational costs. Some recent studies even indicate that this extra computation might itself enhance model capabilities (Pfau et al., 2024). In addition, adaptive compute allocation during inference has been explored. For example, Goyal et al. (2024) incorporated pause tokens into the vocabulary, allowing models to distribute compute more efficiently and improve both reasoning and overall task performance. LightTransfer (Zhang et al., 2024c) introduces a lightweight method that detects lazy layers and replaces their full attention with streaming attention—slashing KV-cache overhead and boosting throughput.
@@ 2.3 Alternatives to Transformer Architectures
Even though most reasoning models are based on the Transformer architecture (Grattafiori et al., 2024; Qwen et al., 2025), alternatives have been proposed to alleviate their high computational cost. Models built on top of RNNs (Beck et al., 2024; Peng et al., 2023), state space models (SSMs) (Gu et al., 2022; Gu & Dao, 2024), and linear attention mechanisms (Katharopoulos et al., 2020; Yang et al., 2024) demonstrate superior inference and memory efficiency, particularly for long-context tasks and large-batch generation. The Mamba series (Mamba-1 and Mamba-2) notably introduced selective state spaces to enable linear-time sequence modeling with strong performance (Gu & Dao, 2024; Dao & Gu, 2024). In addition, hybrid architectures that combine a few self-attention layers with subquadratic layers (e.g., Mamba) have emerged, showing advantages over both pure Transformer and pure subquadratic designs (Lieber et al., 2024; Ren et al., 2024). Such architectures are particularly suited to meet the high compute demands of inference-time scaling, and our work investigates their scaling properties.
@@ 2.4 Knowledge Distillation Strategies
Knowledge distillation has proven to be an effective means of transferring capabilities from large teacher models to smaller, more efficient student models (Hinton et al., 2015). In LLMs, this process compresses a larger pre-trained model into a more compact version while preserving core knowledge and functionality (Gu et al., 2024; Xu et al., 2024). Although larger models tend to exhibit superior reasoning abilities due to scaling properties (Xu et al., 2025; Wei et al., 2022), distillation techniques have enabled smaller models to achieve competitive reasoning performance (DeepSeek-AI et al., 2025; Labs, 2025). While most efforts have focused on intra-architecture distillation (e.g., Transformer-to-Transformer), recent studies have ventured into cross-architecture distillation. For instance, pretrained Transformers have been distilled into architectures such as RNNs (Kasai et al., 2021; Mercat et al., 2024), linear attention models (Zhang et al., 2024b; ), convolutional networks (Ralambomihanta et al., 2024), and SSMs (Bick et al., 2024; Wang et al., 2024b; Paliotta et al., 2025). Whether the robust reasoning abilities of Deepseek R1 (DeepSeek-AI et al., 2025) distilled models can be effectively transferred across different architectures remains an open question.
@@ 3 The M1 Reasoning Model
In this section, we present a multi-stage process for building our hybrid linear RNN reasoning model, M1. The approach has three stages: distillation, SFT, and RL. We begin by distilling a Transformer model into a Mamba architecture, adapting the method of Wang et al. (2024a), which initializes the hybrid model’s weights from a transformer model. We then perform math-specific supervised fine-tuning (SFT) on general mathematical datasets to enhance the model’s mathematical performance, first without yet incorporating datasets generated by reasoning-focused models, and then with reasoning data leveraging multiple large-scale datasets generated by the R1 model series. Finally, we apply R1’s GRPO method to further enhance the model’s math reasoning capability.
@@
The first step in building our M1 model is distilling a pretrained transformer model into a Mamba model. We adapt the distillation approach introduced by  Wang et al. (2024a).
1:Shapes: $B$ - Batch, $L$ - Length, $D$ - embed size, $D/\text{Attention\_heads},NN^{\prime} -$ expand
@@
[…]
@@
The MambaInLlama framework (Wang et al., 2024a) proposes distilling hybrid Transformer-Mamba models by reusing weights from attention layers. In this distillation procedure, outlined in Algorithm 1, linear projections for $\mathbf{Q}$}, $\mathbf{K}$}, $\mathbf{V}$}, and $\mathbf{O}$} are initialized from the corresponding projections for $\mathbf{C}$}, $\mathbf{B}$}, $\mathbf{X}$}, and $\mathbf{O}$}, respectively. The newly introduced parameters in the Mamba layers are the sampling rate $\Delta$ and the dynamic parameter $\mathbf{A}$}, which control the resulting Mamba module via a discretization function. Specifically, the sampling rate $\Delta\in\mathbb{R}^{N^{\prime}}$ discretizes ${\mathbf{B}}_{t},{\mathbf{C}}_{t}\in\mathbb{R}^{N\times 1}$, yielding $\overline{{\mathbf{B}}}_{t},\overline{{\mathbf{C}}}_{t}\in\mathbb{R}^{N^{\prime}\times N\times 1}$, as detailed in Algorithm 1. Different from Wang et al. (2024a), we introduce two additional linear layers to project from head_dim * kv_head to head_dim * n_head. This is because GQA  (Ainslie et al., 2023) is used in the transformer model to reduce the KV cache. As Mamba does not utilize a KV cache, this expansion can increase the expressiveness of $\mathbf{B}$ and $\mathbf{X}$.
@@
After the initial fine-tuning stage, we finetune on an additional set of math problems and solutions generated by reasoning models. We collect a mixed reasoning dataset, including OpenR1-Math-220k  [^2] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k, OpenThoughts-114k-math [^3] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k, and ServiceNow-AI-R1-Distill [^4] https://huggingface.co/datasets/ServiceNow-AI/R1-Distill-SFT, Magpie-Reasoning-250K [^5] https://huggingface.co/datasets/Magpie-Align/Magpie-Reasoning-V2-250K-CoT-Deepseek-R1-Llama-70B for a total of 10B reasoning tokens. The first two datasets were generated from R1, while the last two was generated from the R1 distilled Qwen 32B model and R1 distilled Llama 70B model. We extended the training length to 24,576 because we found that it covers 99% of the data items. We train the model for five epochs using the same optimizer as before but changing the peak learning rate to $6\times 10^{-6}$.
@@ 4 Experiments
$\text{pass@k}=\frac{1}{\text{\# of problems}}\sum_{i=1}^{\text{\# of problems}}\left(1-\frac{\binom{N-C_{i}}{k}}{\binom{N}{k}}\right)$
@@ 4.1 Reasoning Evaluation
Model AIME25 AIME24 MATH500 AMC23 OlympiadBench
Qwen2.5-Math-7B-Instruct - 13.3 79.8 50.6 40.7
rStar-Math-7B (Guan et al., 2025) - 26.7 78.4 47.5 47.1
Eurus-2-7B-PRIME (Cui et al., 2025) - 26.7 79.2 57.8 42.1
Qwen2.5-7B-SimpleRL (Zeng et al., 2025) - 26.7 82.4 62.5 43.3
DeepSeek-R1-Qwen-1.5B 23.0 28.8 82.8 62.9 43.3
M1-3B 23.5 28.9 82.1 62.8 47.3
Table 1: Evaluation results for M1-3B, DeepSeek-R1-Distill-Qwen-1.5B and other MATH models on MATH benchmarks
Model AIME25 AIME24 MATH500 AMC23 OlympiadBench
Pass@1 Maj@32 Pass@1 Maj@32 Pass@1 Maj@32 Pass@1 Maj@32 Pass@1 Maj@32
DeepSeek-R1-Qwen-1.5B 23.0 35.0 28.8 49.2 82.8 91.0 62.9 54.2 43.3 80.3
M1-3B 23.5 34.6 29.0 50.5 82.1 91.8 62.8 55.0 47.3 80.1
Table 2: Maj@32 results comparing M1-3B with DeepSeek-R1-Distill-Qwen-1.5B.
@@
We report the accuracy of M1-3B and DeepSeek-R1-Distill-Qwen-1.5B in Table 1 and  2. We use the baseline DeepSeek-R1-Distill-Qwen-1.5B since a 3B R1 reasoning model is still not available. Although M1-3B has more parameters than DeepSeek-R1-Distill-Qwen-1.5B, its speed is still comparable even with shorter contexts, so we believe this is a fair comparison. Our model’s performance is competitive with state-of-the-art open reasoning models in the same model size range and outperforms larger nonreasoning math transformer models. Our model performs slightly worse on AIME24 compared to the DeepSeek-R1-Distill-Qwen-1.5B model. Notably, DeepSeek-R1-Distill-Qwen-1.5B is built on top of the Qwen2.5 MATH models, which were finetuned with over 1T MATH tokens on top of the Qwen2.5 models, significantly more training data than what M1-3B used in total.
@@ 4.3 Test-Time Scaling
The left side of Figure 3 shows the effect of scaling the number of generated samples (while fixing the maximum decoding length) on AIME25 accuracy. Both the baseline model and M1 see increasing accuracy as the number of samples increases, with M1 nearly matching the baseline performance for larger sample sizes. The efficient generation of M1 also means that generating large number of samples at test-time is faster than for the baseline transformer model.
@@ 5 Analysis
Distill + SFT(MATH) + SFT(Reason) 74 22
Distill + SFT(MATH) + SFT(Reason) + RL 82 28
@@
With more efficient models, we can increase the length of sequences used in RL training, resulting in improved performance. Empirically, we see this in Figure 5, which shows an increase in accuracy on AIME25 as we scale up the length of sequences generated when training with GRPO. Training with sequences of maximum length 4096 results in accuracy below 10%, while allowing sequences up to length 24K boosts the accuracy up to 23%.
@@
Direct Distillation from Reasoning Models We also attempted to distill from Deepseek-R1-Qwen-1.5B instead of Llama-3.2-3B. In this case, we did not SFT on OpenMathInstruct, and instead only SFT on the 10B reasoning data that we collected after distillation. We found that the distilled model’s performance was poor (38% and 3.3% pass@1 accuracy on MATH500 and AIME24, resspectively). Our hypothesis for why this occurs is that 10B tokens is insufficient to effectively transfer reasoning skills from the transformer to Mamba. Although curating a high-quality reasoning dataset demands significant time and effort, we begin by leveraging the standard MATH distillation dataset from OpenMathInstruct (Toshniwal et al., 2024) to first distill a strong MATH model. We then transform this MATH model into a reasoning model via SFT on the dedicated reasoning dataset. This approach achieves strong performance with a much smaller number of reasoning tokens.